OpenStudy (anonymous):

5/ 4 x+1/ 8 x=9/ 8 +x the slashes mean its a fraction i need the steps to solving it also cause i want to understand

jimthompson5910 (jim_thompson5910):

What's the LCD?

OpenStudy (anonymous):

i do not understand the question. is the left hand side \(\frac{9}{8+x}\) ?

OpenStudy (anonymous):

i mean right hand side?

OpenStudy (anonymous):

no its\[\frac{ 5 }{ 4 }x +\frac{ 1 }{ 8 }x=\frac{ 9 }{ 8 }+x\]

jimthompson5910 (jim_thompson5910):

Do you know how to find the LCD?

OpenStudy (anonymous):

\[\frac{ 11 }{ 8}x=\frac{ 9 }{ 8 }+x\]

jimthompson5910 (jim_thompson5910):

now subtract x from both sides

OpenStudy (anonymous):